For the incline shown below, how many ft in the vertical direction correspond to 4ft in the horizontal direction? ( h=12 v=3)
Answer by Alan3354(69443)

Slope m = v/h = 3/12 = 1/4
------------------
1/4 = h/4
h = 1 ft